Frank drives his car 380 miles and has a mean of a certain speed. If the mean speed had been 2mph more, he could have traveled 399 miles in the same length of time. What was his mean speed?

Answer by Boreal(15235)

mean*time=distance; let mean=x
first time=380/x
second time=399/(x+2)
but the times are equal
therefore, (380/x)=(399/(x+2))
cross-multiply
380(x+2)=399x
380x+760=399x
19x=760
x=40 mph ANSWER
It took 9.5 hours to go 380 miles
at 42 mph and 9.5 hours he goes 399 miles.